An inventor has developed a new, energy efficient lawn mower engine. He claims that the engine will run continuously for 5 hours (300 minutes) on a single gallon of regular gasoline. Suppose a simple random sample of 50 engines is tested. The engines run for an average of 295 minutes, with a standard deviation of 20 minutes. Test the hypothesis that the mean run time is 300minutes against the alternative hypothesis that the mean run time is not 300minutes using a 0.05 level of significance


1
Expert's answer
2022-02-28T11:18:28-0500

"H_0:\mu=300."

"H_a:\mu \ne300."

Test statistic: "t=\frac{\bar x-\mu}{\frac{s}{\sqrt{n}}}=\frac{295-300}{\frac{20}{\sqrt{50}}}=-1.77."

Degrees of freedom: "df=n-1=50-1=49."

P-value: "p=2P(T<-1.77)=0.0833."

Since the p-value is greater than 0.05, fail to reject the null hypothesis.

There is no sufficient evidence that the mean run time is not 300 minutes.